The story so far
This presentation is the property of its rightful owner.
Sponsored Links
1 / 9

The Story So Far PowerPoint PPT Presentation


  • 93 Views
  • Uploaded on
  • Presentation posted in: General

The Story So Far. Cold in both rings. 4K Wave. PP12 lattice test evening. PS work during Day. Overnight Stores 109x109. PS gave us e-lens. First Collisions. Start Physics. LIPA Maintenance. 6 hour stores with tune between. LIPA Maintenance. 8hr Stores. APEX. Inject =200e11.

Download Presentation

The Story So Far

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


The story so far

The Story So Far

Cold in

both rings

4K Wave

PP12 lattice test evening

PS work during Day

Overnight Stores

109x109

PS gave

us e-lens

First Collisions

Start Physics


The story so far

LIPA

Maintenance

6 hour stores with tune between

LIPA

Maintenance

8hr Stores

APEX

Inject =200e11


The story so far

Maintenance

Switch Fy12

Lattice

Pushed

Intensity

Reach 200e11

At FT

Rotator Tuning

APEX

Maintenance

MD Studies

Golden Store 17396

Pushed Bunch

Intensity

APEX

Machine Dev.

APEX

Machine Dev.

Test corrected

Optics


The story so far

MD Studies

Installed Dpx

APEX

Maint.

Dpx reduction

In Blue removed

MD

Spin Tune

Low E APEX


Past week

Past week

  • Machine Development

    • RF group Damper studies

    • Removed Dpx reduction in Blue

  • Trouble with aborts

  • Steering Star back into full collision

  • Angles at IP's taken out


Current situation

Current Situation

  • As compared to Period (17416-17451) Blue Jet average = 58% we are still low in Blue

    • During Period with Dpx reduction applied in Blue average Jet numbers for Blue = 51%

    • Since removal of Dpx reduction Blue Jet = 53%

    • Not clear that the Dpx reduction was a culprit, however Emittances on the ramp have seemed to blow up on the ramp since 17545 after removal of Dpx reduction so this may obscure issue. Also some of the fills had lower AGS injection numbers. Need more analysis.

  • Yellow Appears 'healthy' any low jet numbers appear to due to low AGS and/or source numbers.


Blue emittance blow up on energy ramp

Blue emittance blow up on energy ramp


Blue shorter bunch length

Blue shorter bunch length


Future plans

Future Plans

  • Run low intensity fill to help diagnose abort situation

  • Sort out why angle's creeping back into IP's

  • Spin Tune measurement at FT without rotators to make sure snake currents correct

  • RF bb damper work to improve performance

  • Better understand causes of emittance growth in Blue on energy ramp.


  • Login